A 25-kg chandelier hangs from a ceiling on a vertical 4.0-m-long wire.

Part A.

What horizontal force would be necessary to displace its position 0.12 m to one side?

Part B

What will be the tension in the wire?
Answer #1

A)

at 12 m :

Angle from vertical, x= asin (0.12/4) = 1.72 degree

At this point:

T cos x = m*g

T*cos 1.72 = 25*9.8

T= 245.11 N

Forceapplied must baalnce T* sin x

T sin x = F

F = 245.11 * sin 1.72

F=7.36 N

B)

As calculated above,Tenstion, T= 245.11 N